Dont Fall Down Incorporated produces skates and offers a one-year warranty on all products. During 2019, the company recorded net sales of $3,887.4 million. Historically, about 2% of all sales are returned under warranty and the cost of repairing and or replacing goods under warranty is about 50% of retail value. Assume that at the start of the year Dont Fall Down’s balance sheet included an accrued warranty liability of $16.9 million and at the end of the year, the accrued warranty liability balance was $13.0 million. (Hint: A T-account of Warranty Liability may help you in answering this question). (Hint 2: This is a liability account. Determine what would make it increase and what would decrease the Warranty Liability account).

Calculate Dont Fall Down’s warranty expense for 2019. How much did Dont Fall Down pay during the year to repair and or replace goods under warranty?
Answer #1

Warranty expense = 3887.4*2% = $77.748 million

b) Amount paid = Brginning liability + accrued during the year - balance at end

= 16.9+77.748-13

= $81.648 million